# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v2

@@ Abstract
In interactive decision-making tasks, information can be acquired by direct interactions, through receiving indirect feedback, and from external knowledgeable sources. We examine the trade-off between the information an agent accumulates and the regret it suffers. We show that information from external sources, measured in bits, can be traded off for regret, measured in reward. We invoke information-theoretic methods for obtaining regret lower bounds, that also allow us to easily re-derive several known lower bounds. We then generalize a variety of interactive decision-making tasks with external information to a new setting. Using this setting, we introduce the first Bayesian regret lower bounds that depend on the information an agent accumulates. These lower bounds also prove the near-optimality of Thompson sampling for Bayesian problems. Finally, we demonstrate the utility of these bounds in improving the performance of a question-answering task with large language models, allowing us to obtain valuable insights.
@@ 1 Introduction
In interactive decision-making tasks, an agent interacts with an environment by sequentially choosing decisions from a decision space. Subsequently, the agent receives feedback that usually includes a reward and, optionally, other types of information. For example, in multi-armed bandits (MAB) [Lattimore and Szepesvári, 2020], the feedback includes only the reward. In partial monitoring, however, it only includes a signal, which the agent then needs to use to indirectly infer the reward [Cesa-Bianchi et al., 2006]. The goal of the agent in such tasks is to minimize the gap between the accumulated reward and the reward of the best decision in hindsight, also known as regret.
In these tasks, agents can accumulate information from multiple sources, that can be categorized into two main types. One is information acquired through direct interactions and receiving feedback from the task. For example, in reinforcement learning such information includes visited states, accumulated rewards, and performed actions. The other type is information provided by external sources. Such information includes human advice [Najar and Chetouani, 2021], text description generated by large language models [Du et al., 2023, Hu and Sadigh, 2023], and more. Figure 1 describes a general decision-making task with external information. Resource allocation is an example of a task that combines both types of information. In this task, our goal is to route a process from a mobile device to a server while minimizing energy consumption and execution time [Zuo and Joe-Wong, 2021]. In this task, we can accumulate information directly by routing processes and receiving feedback. However, we also have relevant information that can be obtained from other sources, such as server specifications, load information, and more.
When we receive information from multiple sources, we can measure the total information in bits. However, we assess the algorithm’s performance using regret. Thus, a question arises:
@@
Figure 1: Schematic of general interactive decision-making with contextual information task. Every round, a source of knowledge provides our agent with information. The agent then makes a decision, that causes the task to generate an observation and a reward. The observation is revealed to the agent, who updates his next decision according to the past observations and the information he received.
@@ 1.1 Contributions
We introduce a novel information-theoretic method to obtain regret lower bounds for a wide variety of interactive decision-making problems. This method uses Fano’s inequality [Cover, 1999] and the construction method introduced by Yang and Barron [1999] to lower bound the worst-case Bayesian regret. We then use this method to derive lower bounds for multi-armed bandits, contextual bandits, and reinforcement learning tasks. This method allows us to lower-bound both the Bayesian and frequentist regret.
Information-theoretic prior-dependent lower bounds
can be obtained using this method. We consider an interactive decision-making framework that describes general online tasks with contextual information in a Bayesian setting. We examine the regret of policies with a mutual information constraint and obtain Bayesian regret lower bounds that depend on the entropy of our prior. These bounds also show that both Thompson sampling [Russo and Van Roy, 2016] and information-directed sampling (IDS) [Russo and Van Roy, 2014] are nearly optimal in the Bayesian sense for multi-armed and linear bandits.
@@
We use the lower bounds to quantify the relationship between information, measured in bits, and regret. We show that the presented lower bound, combined with existing upper bounds, quantifies the relationship between external information and regret. Furthermore, these bounds quantify the relationship between information accumulation and regret in online tasks. Finally, we utilize the trade-off in a question-answering task with large language models, demonstrating how we can easily utilize this trade-off in online settings.
@@ 2 Setting and Preliminaries
We introduce a new setting called interactive decision-making in Bayesian environments, with contextual information. This setting is a Bayesian adaptation of the frequentist interactive decision-making setting introduced by Foster et al. [2021]. We denote the set of Borel probability measures over a locally compact space by $\Delta(\cdot)$. Let $\Pi$ be a compact decision space and $\mathcal{M}$ be a compact class of models, where each model $M\in\mathcal{M}$ is a mapping from the decision space to an observation space $\mathcal{O}$, i.e., $M:\Pi\to\mathcal{O}$. Furthermore, every model $M$ has a reward function, $R_{M}:\Pi\to\Delta([0,1])$ that maps every decision to a reward distribution over $[0,1]$. $\mathcal{C}$ is another compact space called the context space. A task in this setting is defined by the decision space $\Pi$, the model class $\mathcal{M}$, the context space $\mathcal{C}$, and the prior $P\in\Delta(\mathcal{M}\otimes\mathcal{C})$. We denote by $\mu_{M}(\pi)=\mathbb{E}_{x\sim R_{M}(\pi)}[x]$ the mean reward of decision $\pi$ under model $M$. With a slight abuse of notation, we also denote by $\mu_{M}(\phi)=\mathbb{E}_{\pi\sim\phi}[\mu_{M}(\pi)]$, the mean reward of a policy $\phi\in\Delta(\Pi)$ under model $M$.
@@
In a $T$-round game, a random model and context $(M_{t},C_{t})$ are sampled every round, independently from previous rounds, according to the prior probability $P$. Before each round starts, the context $C_{t}$ is revealed to the agent, who determines his policy $p_{t}\in\Delta(\Pi)$ according to the context and the observed history. Then, a decision $\pi_{t}\sim p_{t}$ is sampled and an observation $O_{t}\sim M_{t}(\pi_{t})$ is revealed to the agent. A reward, $r_{t}\sim R_{M_{t}}(\pi_{t})$ is also generated. We denote the history up to time $t$ as $\mathcal{H}_{t}=\{C_{i},\pi_{i},O_{i}\}_{i=1}^{t}$. We measure the performance with respect to the best context-dependent decision in hindsight. This performance metric is called the Bayesian regret and is defined by
$\mathcal{BR}_{P}(T;\{\pi_{t}\}_{t=1}^{T})=\sup_{\pi^{*}:\mathcal{C}\to\Pi}\mathbb{E}\left[\sum_{t=1}^{T}\left(R_{M_{t}}(\pi^{*}({c_{t}}))-R_{M_{t}}(\pi_{t})\right)\right]$
@@
where the expectation is taken with respect to the randomness in decisions, observations, and rewards ($(M_{t},C_{t})\sim P$, $\pi_{t}\sim p_{t}$ and $r_{t}\sim R_{M_{t}}(\pi_{t})$). We also note that by fixing a single context, $\left|\mathcal{C}\right|=1$, the setting reduces to regular Bayesian interactive decision-making. When obvious from the context, $P$ and $\{\pi_{t}\}_{t=1}^{T}$ are omitted. We now show how our setting covers common decision-making tasks with contextual information.
Example 2.1 (Contextual multi-armed bandit with Bernoulli rewards).
@@
We demonstrate how our general framework reduces to the contextual MAB problem, where the rewards for each arm are Bernoulli random variables. Assume a finite set of decisions, which are called arms in this setting, $\Pi=\{1,2,\dots,A\}$. Furthermore, assume that the context space is finite as well, $\mathcal{C}=\{1,2,\dots,C\}$. The model class is then given by =$\mathcal{M}=\{f|f:\Pi\times\mathcal{C}\to[0,1]\}$, so every model maps $\left(\pi,C\right)$ to $\mathrm{Bern}(f(\pi,C))$. At each round $t$, a model and context are sampled $(M_{t},C_{t})\sim P$, and the context is revealed. A decision $\pi_{t}$ is selected and the observation is the incurred reward, $r_{t}=O_{t}\sim\mathrm{Bern}(f_{t}(\pi_{t},C_{t}))$.
@@
We also demonstrate how our setting applies to a finite horizon MDP. We consider a tabular MDP with a finite horizon, which is defined according to the tuple $(\mathcal{S},\mathcal{A},\mathbb{P},\mathbb{R},H)$. We also assume the episodic and stationary setting, where we update the policy only at the beginning of every episode. Additionally, we assume that the reward of every episode is bounded in $[0,1]$. In this setting, the decision space is all of the deterministic policies, so $\Pi=\{\pi:\mathcal{S}\times\mathcal{H}\to\mathcal{A}\}$. The model class $\mathcal{M}$ is defined by the set of all MDPs with the same state space, action set, and horizon. At every time step, the agent determines a policy and plays it over the entire horizon. The observation is the trajectories, which include the visited states, rewards, and actions performed.
Covering and packing
@@
Assume a normed space $A$ with metric $d$ induced by the norm, and let $\epsilon>0$. A set $F\subseteq A$ is called a covering set of $A$ if for all $a\in A$, there exists $f\in F$ such that $d(a,f)\leq\epsilon$. The $\epsilon$-covering number of $A$, denoted by $\mathcal{N}(d,A,\epsilon)$, is the smallest cardinality possible for an $\epsilon$-covering set. A set $G\subseteq A$ is called a packing set of $A$ if $\forall g_{1},g_{2}\in G$, $d(g_{1},g_{2})\geq\epsilon$. The $\epsilon$-packing number of $A$ is the largest possible cardinality for an $\epsilon$-packing set. We define an $\epsilon$-local set of A as $B(b,\epsilon)=\{a\in A:d(b,a)\leq\epsilon\}$, for every $b\in A$ and $\epsilon>0$. The local packing number, denoted by $\mathcal{M}^{\mathrm{local}}\left(d,A,\epsilon\right)$, is the largest $(\epsilon/2)$-packing set of any set $B(b,\epsilon)$, i.e.,
@@
$\mathcal{M}^{\mathrm{local}}\left(d,A,\epsilon\right)=\max\left\{M:\exists b\text{ such that the }(\epsilon/2)\text{-packing number of }B(b,\epsilon)\text{ is }M\right\}.$
@@
A set $E\subseteq A$ will be called an $\epsilon$-local packing set if $\exists b$ such that $E$ is an $(\epsilon/2)$-packing set of $B(b,\epsilon)$ and $\left|E\right|=\mathcal{M}^{\mathrm{local}}\left(d,A,\epsilon\right)$. When $d(x,y)=\left\|x-y\right\|_{p}$ we denote $\mathcal{M}_{p}(A,\epsilon),\mathcal{N}_{p}(A,\epsilon)$ and $\mathcal{M}^{\mathrm{local}}_{p}\left(A,\epsilon\right)$ for the $\epsilon$-packing, $\epsilon$-covering and $\epsilon$-local packing number, respectively.
Additional notations
@@
We denote by $H(X)=\mathbb{E}\left[-\log Q(X)\right]$ the Shannon entropy of a random variable $X\sim Q$, and with a slight abuse of notation also denote $H(Q)=H(X)$. Given two jointly distributed random variables, $(X,Y)\sim Q$, we denote their marginal distributions by $Q(x)$ and $Q(y)$. We also define the mutual information of two jointly distributed random variables $(X,Y)\sim Q$ as $I(X;Y)=\mathbb{E}\left[\log\left(\frac{Q(X,Y)}{Q(X)Q(Y)}\right)\right]$. The KL-divergence is defined as $D\left(P,Q\right)=\mathbb{E}_{X\sim P}\log\frac{P(X)}{Q(X)}$ Unless stated otherwise, logarithms are assumed to be in base-2, denoted by $\log$. We also denote $x\gtrsim y$ if there is some global constant $c>0$ such that $x\geq cy$.
Table 1: Lower bounds for MAB and RL problems, recovered with Theorem 3.5 and Proposition 4.3.
@@
MAB $\Omega(\sqrt{KT})$ Yes
Tabular, episodic and stationary MDP $\Omega(\sqrt{HSKT})$ Yes [Zhang et al., 2021]
@@
Bayesian MAB ${\Omega\left(\sqrt{AT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)}$ First bound
Bayesian contextual MAB ${\Omega\left(\sqrt{CAT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)}$ First bound
Bayesian linear bandit $\Omega\left(\sqrt{dT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$ First bound
@@ 3 Regret Lower Bounds Using Fano’s Inequality
$\inf_{\{\pi_{t}\}_{t=1}^{T}}\sup_{\nu\in\Delta(\mathcal{M}\times\mathcal{C})}\mathcal{BR}_{\nu}(T;\{\pi_{t}\}_{t=1}^{T})\equiv\mathcal{BR}^{*}(T).$
@@
Theorem 3.1 (Fano’s inequality, Theorem 2.10 of [Cover, 1999]).
Let $X,Y\sim Q$ be two jointly distributed random variables, where $X$ can take values over a finite set, whose cardinality is $\mathcal{X}$. Let $\hat{X}=f(Y)$ for some $f$ be an estimator of $X$. Then the following holds for all $f$,
@@
Fano’s inequality can be used to lower-bound the error in hypothesis testing problems. Thus, to utilize Fano’s inequality, we need to reduce our problem from regret minimization to hypothesis testing. We assume access to some set $\Phi=\{\phi_{1},\dots,\phi_{K}\}$ of (possibly stochastic) decisions such that for some $\epsilon>0$, $\mathbb{E}_{M\sim P}[\left|\mu_{M}(\phi_{i})-\mu_{M}(\phi_{j})\right|]\geq\epsilon$, for all $i\neq j$. From Markov’s inequality, we have that for any $\phi\in\Delta(\Pi)$,
$\sup_{i}\mathbb{E}_{M\sim P}[\left|\mu_{M}(\phi_{i})-\mu_{M}(\phi)\right|]\geq\epsilon\sup_{i}\mathbb{P}\left(\left|\mu_{M}(\phi_{i})-\mu_{M}(\phi)\right|\geq\epsilon\right).$ (2)
Now, instead of lower-bounding the regret, we lower-bound the error probability. This reduces our problem from regret minimization to hypothesis testing. We also assume that a metric $\rho$ exists such that $\mathbb{E}[\left|\mu_{M}(\psi)-\mu_{M}(\phi)\right|]\geq\rho(\phi,\psi)$ for all $\psi,\phi\in\Delta(\Pi)$. Using Theorem 3.1, we can lower-bound the prediction error. This results in the following proposition.
@@
For any algorithm, prior $P$, a decision set $\Phi_{K}=\{\phi_{1},\dots,\phi_{K}\}\subseteq\Delta(\Pi)$ and for all $\epsilon>0$ such that $\rho(\phi_{i},\phi_{j})\geq\epsilon$ for all $i\neq j$,
$\frac{\mathcal{BR}_{P}(T)}{T\epsilon}\geq 1-\frac{I(V;\mathcal{H}_{T})+1}{\log K}$ (3)
@@
Proposition 3.2 allows us to lower-bound the Bayesian regret using a specific set of decisions. We suggest a method for selecting the decision set $\Phi$, which provides a general approach for obtaining regret lower bounds. We divide our analysis into finite, parametric with infinite decisions, and non-parametric decision spaces. To make the lower bound in Equation 3 tighter, we should select the largest value of $K$ that satisfies the requirements for a given $\epsilon>0$. Thus, for finite decision spaces, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision simplex, $\Delta(\Pi)$, according to $\rho$, under $P$. Thus, we have $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$. Similarly, for parametric decision spaces with infinite decisions, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision set, $\Pi$, i.e., $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Pi,\epsilon)$. Yang and Barron [1999] have obtained a relationship between the local and global packing numbers, which we state in Lemma 3.3.
Lemma 3.3 (Lemma 3 of Yang and Barron [1999]).
For any space $A$ with metric $d$ and for all $\epsilon>0$,
$\log{\mathcal{M}}\left(d,A,\frac{\epsilon}{2}\right)-\log{\mathcal{M}}(d,A,\epsilon)\leq\log{\mathcal{M}}^{\mathrm{loc}}(d,A,\epsilon).$
For non-parametric decision spaces, we use an $\epsilon$-global packing set of our decision space $\Pi$, which means that $K=\mathcal{M}(\rho,\Pi,\epsilon)$. After selecting $\Phi$, we need to handle $I(V;\mathcal{H}_{T})$. This can be achieved using the following theorem.
@@
Under the assumptions and notations of Proposition 3.2, if $\Pi$ is a parametric decision space then
$I(V;\mathcal{H}_{T})\lesssim 2\epsilon^{2}T.$
If $\Pi$ is a non-parametric decision space then
$I(V;\mathcal{H}_{T})\leq\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right).$
Theorem 3.4 allows us to obtain upper bounds for the mutual information. Furthermore, it provides us with a relationship between the global and local packing numbers. Combining Theorem 3.4, Proposition 3.2 and the selection of $\Phi$, results in the following Proposition.
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T+1}{\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon/2)-\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon)}.$ (4)

### v3

@@ Abstract
In many sequential decision problems, an agent performs a repeated task. He then suffers regret and obtains information that he may use in the following rounds. However, sometimes the agent may also obtain information and avoid suffering regret by querying external sources. We study the trade-off between the information an agent accumulates and the regret it suffers. We invoke information-theoretic methods for obtaining regret lower bounds, that also allow us to easily re-derive several known lower bounds. We introduce the first Bayesian regret lower bounds that depend on the information an agent accumulates. We also prove regret upper bounds using the amount of information the agent accumulates. These bounds show that information measured in bits, can be traded off for regret, measured in reward. Finally, we demonstrate the utility of these bounds in improving the performance of a question-answering task with large language models, allowing us to obtain valuable insights.
@@ 1 Introduction
In interactive decision-making problems, an agent repeatedly interacts with a task by sequentially choosing decisions from a decision space. Subsequently, the agent receives feedback that usually includes a reward and, optionally, other types of information. For example, the feedback includes only the reward in multi-armed bandit (MAB) problems [Lattimore and Szepesvári, 2020]. In partial monitoring, however, it only includes a signal, which the agent then utilizes to indirectly infer the reward [Cesa-Bianchi et al., 2006]. The goal of the agent in such tasks is to minimize the gap between the accumulated reward and the reward of the best decision in hindsight, also known as regret.
In these tasks, agents can accumulate information from multiple sources, that can be categorized into two main types. One is information acquired through direct interactions and receiving feedback from the task. For example, in reinforcement learning (RL) such information includes visited states, accumulated rewards, and performed actions. The other type is information provided by external sources, such as human advice [Najar and Chetouani, 2021], text description generated by large language models [Du et al., 2023, Hu and Sadigh, 2023], and more.
Resource allocation is an example of a task that combines both types of information. In this task, our goal is to route a process from a mobile device to a server while minimizing energy consumption and execution time [Zuo and Joe-Wong, 2021]. In this task, we can accumulate information directly by routing processes and receiving feedback. However, we also have relevant information that can be obtained from other sources, such as server specifications, load information, and more.
Consider two agents playing on the same online task. Both agents have access to the same observations, but the first agent has already played multiple rounds in the task. Thus, he already suffered regret and gained an advantage over the second agent. Before he starts interacting with the task, the second agent can use external knowledge that will allow him to reach the same performance level as the first agent. In this paper, we focus on how much information the second agent needs to query so he will "catch up" with the first agent without suffering regret. In other words, we focus on the following question:
@@
Figure 1: Schematic of a general interactive decision-making task with contextual information. Every round, a source of knowledge provides our agent with information. The agent then makes a decision, that causes the task to generate an observation and a reward. The observation is revealed to the agent, who updates his next decision according to the past observations and the information he received.
@@ 1.1 Contributions
We introduce a novel information-theoretic method to obtain regret lower bounds for a wide variety of interactive decision-making problems. This method uses Fano’s inequality [Cover, 1999] and the construction method introduced by Yang and Barron [1999] to lower bound the worst-case Bayesian regret. We then use this method to derive lower bounds for MAB, contextual bandit, and RL tasks.
Information-theoretic prior-dependent bounds.
We consider an interactive decision-making framework that describes general online tasks with contextual information in a Bayesian setting. We examine the regret of policies with a mutual information constraint and obtain Bayesian regret lower bounds that depend on the amount of information an agent accumulates. Additionally, we show upper bounds for Thompson sampling [Thompson, 1933] in an MAB environment, which depends on the information the agent collects. We also present lower bounds for scenarios where the entropy of the prior is constrained, for a specific family of algorithms.
@@
We show that the presented lower bounds quantify the relationship between external information and regret. Furthermore, these bounds quantify the relationship between the rate of information accumulation and regret in online tasks. These relationships allow us to measure how much regret can be avoided by looking at the information that can be accumulated. We then demonstrate how to easily utilize this insight in an online question-answering task with large language models.
The rest of the paper is organized as follows: In Section 2 we present our setting and relevant preliminaries. In Section 3 we present a novel information-theoretic method for obtaining regret lower bounds. In Section 4 we present regret bounds that depend on the information accumulated by the agent, followed by experiments in Section 5. In Section 6 we review related work, and draw conclusions in Section 7.
@@ 2 Setting and Preliminaries
We introduce a new setting called interactive decision-making in Bayesian environments with contextual information, which is illustrated in Figure 1. This setting is a Bayesian adaptation of the frequentist interactive decision-making setting introduced by Foster et al. [2021]. We denote the set of Borel probability measures over a locally compact space by $\Delta(\cdot)$. Let $\Pi$ be a compact decision space, $\mathcal{M}$ be a compact model space and $\mathcal{C}$ be a compact context space. Every model $M\in\mathcal{M}$ is a mapping from the decision and context space to an observation space $\mathcal{O}$, i.e., $M:\Pi\times\mathcal{C}\to\mathcal{O}$. Every model $M$ has a reward function, $R_{M}:\Pi\times\mathcal{C}\to\Delta([0,1])$ that maps every decision-context pair to a reward distribution over $[0,1]$. A task in this setting is defined by the decision space $\Pi$, the model class $\mathcal{M}$, the context space $\mathcal{C}$, the reward functions $R_{M}$, and the prior $P\in\Delta(\mathcal{M}\times\mathcal{C})$. We denote by $\mu_{M}(\pi,C)=\mathbb{E}_{x\sim R_{M}(\pi,C)}[x]$ the mean reward of decision $\pi$ under model $M$ and context $C$. With a slight abuse of notation, we also denote by $\mu_{M}(\phi,C)=\mathbb{E}_{\pi\sim\phi}[\mu_{M}(\pi,C)]$, the mean reward of a policy $\phi\in\Delta(\Pi)$ under model $M$ and context $C$.
@@
In a $T$-round game, a random model $M$ is sampled before the game starts according to the prior $P$. Then a context $C_{t}$ is sampled every round, independently from previous rounds, according to the marginal prior probability $P(\cdot\mid M)$. Before each round starts, the context $C_{t}$ is revealed to the agent, who determines his policy $p_{t}:\mathcal{C}\to\Delta(\Pi)$ according to the context and the observed history. Then, a decision $\pi_{t}\sim p_{t}$ is sampled and an observation $O_{t}\sim M(\pi_{t},C_{t})$ is revealed to the agent. A reward, $r_{t}\sim R_{M}(\pi_{t},C_{t})$ is also generated. We denote the history up to time $t$ as $\mathcal{H}_{t}=\{C_{i},\pi_{i},O_{i}\}_{i=1}^{t}$. We measure the performance with respect to the best context-dependent decision in hindsight. This performance metric is called the Bayesian regret and is defined by
$\mathcal{BR}_{P}(T;\{p_{t}\}_{t=1}^{T})=\sup_{\pi^{*}:\mathcal{C}\to\Pi}\mathbb{E}\left[\sum_{t=1}^{T}\left(\mu_{M}(\pi^{*}({C_{t}}),C_{t})-r_{t}\right)\right]$ (1)
@@
where the expectation is taken with respect to the randomness in decisions, contexts, observations, and rewards ($(M,C_{t})\sim P$, $\pi_{t}\sim p_{t}$ and $r_{t}\sim R_{M}(\pi_{t},C_{t})$). We also note that by fixing a single context, $\left|\mathcal{C}\right|=1$, the setting reduces to regular Bayesian interactive decision-making. When obvious from the context, $P$ and $\{p_{t}\}_{t=1}^{T}$ are omitted. We next show how our general framework covers common decision-making tasks with contextual information.
Example 2.1 (Contextual MAB with Bernoulli rewards).
@@
The contextual MAB problem is defined by the rewards of each arm, which are, say, Bernoulli random variables. For simplicity, we assume a finite set of decisions, which are called arms in this setting, $\Pi=\{1,2,\dots,K\}$. Furthermore, we assume that the context space is finite as well, $\mathcal{C}=\{1,2,\dots,C\}$. The model class is then given by =$\mathcal{M}=\{f|f:\Pi\times\mathcal{C}\to[0,1]\}$, so every model maps $\left(\pi,C\right)$ to $\mathrm{Bern}(f(\pi,C))$. At each round $t$, a context is sampled and revealed. A decision $\pi_{t}$ is selected and the observation is the incurred reward, $r_{t}=O_{t}\sim\mathrm{Bern}(f_{t}(\pi_{t},C_{t}))$.
@@
The finite-horizon tabular MDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathbb{P},\mathbb{R},H)$ (which are the state space, action space, transition kernel, reward function and horizon, accordingly). We assume the episodic and stationary setting, where we update the policy only at the beginning of every episode. Additionally, we assume that the reward of every episode is bounded in $[0,1]$. In this setting, the decision space is all of the deterministic policies, so $\Pi=\{\pi:\mathcal{S}\times\mathcal{H}\to\mathcal{A}\}$. The model class $\mathcal{M}$ is defined by the set of all MDPs with the same state space, action set, and horizon. At every time step, the agent determines a policy and plays it over the entire horizon. The observations are the trajectories, which include the visited states, rewards, and actions performed.
Covering and packing.
@@
Consider a normed space $A$ with metric $d$ induced by the norm, and let $\epsilon>0$. We denote by $\mathcal{N}(d,A,\epsilon)$ and $\mathcal{M}(d,A,\epsilon)$ the $\epsilon$-covering and $\epsilon$-packing numbers of $A$ under the norm $d$, respectively. We define an $\epsilon$-local set of $A$ as $B(b,\epsilon)=\{a\in A:d(b,a)\leq\epsilon\}$, for every $b\in A$ and $\epsilon>0$. The local packing number, denoted by $\mathcal{M}^{\mathrm{loc}}\left(d,A,\epsilon\right)$, is the largest $(\epsilon/2)$-packing set of any set $B(b,\epsilon)$, i.e.,
@@
$\mathcal{M}^{\mathrm{loc}}\left(d,A,\epsilon\right)=\max\left\{M:\text{there is }b\text{ such that the }(\epsilon/2)\text{-packing number of }B(b,\epsilon)\text{ is }M\right\}.$
@@
When $d(x,y)=\left\|x-y\right\|_{p}$ we denote by $\mathcal{M}_{p}(A,\epsilon),\mathcal{N}_{p}(A,\epsilon)$ and $\mathcal{M}^{\mathrm{loc}}_{p}\left(A,\epsilon\right)$ the $\epsilon$-packing, $\epsilon$-covering and $\epsilon$-local packing number, respectively. Detailed definitions and properties of the covering, packing and local packing numbers are provided in Appendix B.
Additional notations.
@@
We denote by $H(X)=\mathbb{E}\left[-\log Q(X)\right]$ the Shannon entropy of a random variable $X\sim Q$, and with a slight abuse of notation, we also denote $H(Q)=H(X)$. Given two jointly distributed random variables, $(X,Y)\sim Q$, we denote their marginal distributions by $Q(x)$ and $Q(y)$. We also define the mutual information of two jointly distributed random variables $(X,Y)\sim Q$ as $I(X;Y)=\mathbb{E}\left[\log\left(\frac{Q(X,Y)}{Q(X)Q(Y)}\right)\right]$. The KL-divergence is defined as $D\left(P,Q\right)=\mathbb{E}_{X\sim P}\left[\log\frac{P(X)}{Q(X)}\right]$ [Cover, 1999]. Unless stated otherwise, logarithms are assumed to be in base-2, denoted by $\log$. We also denote $x\gtrsim y$ if there is some global constant $c>0$ such that $x\geq cy$. We define by $\mathbb{B}_{p}^{d}=\{x\in\mathbb{R}^{d}:\left\|x\right\|_{p}\leq 1\}$ the unit sphere in $\mathbb{R}^{d}$ under the $L_{p}$ norm.
Table 1: Lower bounds for online problems, recovered using methods presented in the paper.
@@
MAB $\Omega(\sqrt{KT})$ Yes [Audibert and Bubeck, 2009]
Tabular RL $\Omega(\sqrt{HSKT})$ Yes [Zhang et al., 2021]
@@
Bayesian MAB ($R$ bits) ${\Omega\left(\sqrt{KT\frac{\log K}{R}}\right)}$ First bound
Bayesian linear MAB ($R$ bits) $\Omega\left(\sqrt{dT\frac{\log K}{R}}\right)$ First bound
Bayesian MAB $\Omega\left(\sqrt{KT\frac{H(\pi^{*})}{\log K}}\right)$ First bound
@@ 3 Regret Lower Bounds Using Fano’s Inequality
$\sup_{\nu\in\Delta(\mathcal{M}\times\mathcal{C})}\inf_{\{p_{t}\}_{t=1}^{T}}\mathcal{BR}_{\nu}(T;\{p_{t}\}_{t=1}^{T})\equiv\mathcal{BR}^{*}(T).$ (2)
@@
Theorem 3.1 (Fano’s inequality, Theorem 2.10 of Cover [1999]).
Let $X,Y\sim Q$ be two jointly distributed random variables, where $X$ can take values over a finite set, whose cardinality is $\mathcal{X}$. Let $\hat{X}=f(Y)$ for some $f$ be an estimator of $X$. If $\hat{X}$ is uniformly distributed over all possible values in $\mathcal{X}$, then the following holds for all $f$,
@@
While Theorem 3.1 is used to lower-bound the error in hypothesis testing problems, we now show how it can also be applied to regret minimization. We assume access to some set $\Phi=\{\phi_{1},\dots,\phi_{K}\}$ of (possibly stochastic) policies. We also assume that a metric $\rho$ exists such that $\mathbb{E}[\left|\mu_{M}(\psi,C)-\mu_{M}(\phi,C)\right|]\geq\rho(\phi,\psi)$ for all $\psi,\phi\in\Delta(\Pi)$. The following proposition applies Theorem 3.1 to regret minimization, by reducing it to a hypothesis testing over the finite set $\Phi$.
@@
For any algorithm, prior $P$, a decision set $\Phi=\{\phi_{1},\dots,\phi_{K}\}\subseteq\Delta(\Pi)$ and for all $\epsilon>0$ such that $\rho(\phi_{i},\phi_{j})\geq\epsilon$ for all $i\neq j$,
$\frac{\mathcal{BR}_{P}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{I(V;\mathcal{H}_{T})+1}{\log K}\right]$ (4)
@@
The proof of Proposition 3.2 is provided in Appendix C.1. Proposition 3.2 allows us to lower-bound the Bayesian regret using a specific set of decisions. We suggest a method for selecting the decision set $\Phi$, which provides a general approach for obtaining regret lower bounds. We separate our analysis into finite, parametric with infinite decisions, and non-parametric decision spaces. This separation is done to make the lower bound in Equation 4 tighter. We observe that the lower bound is tighter for larger values of $K$. Thus, we should select the largest value of $K$ that satisfies the requirement $\rho(\phi_{i},\phi_{j})\geq\epsilon$ for a given $\epsilon>0$. For finite decision spaces, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision simplex, $\Delta(\Pi)$, according to $\rho$, under $P$. So, we have $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$. Similarly, for parametric decision spaces with infinite decisions, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision set, $\Pi$, i.e., $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Pi,\epsilon)$. For non-parametric decision spaces, we use an $\epsilon$-global packing set of our decision space $\Pi$, which means that $K=\mathcal{M}(\rho,\Pi,\epsilon)$. After selecting $\Phi$, we need to handle $I(V;\mathcal{H}_{T})$. To do this, we make the following assumption about the policies.
Assumption 3.1.
There exist constants $\bar{A},\epsilon_{0}>0$ such that for all policies $p_{1},p_{2}$ such that $\rho(p_{1},p_{2})\leq\epsilon_{0}$, $D\left(p_{1},p_{2}\right)\leq 2\bar{A}\rho(p_{1},p_{2})^{2}$.
This assumption is also made by Yang and Barron [1999]. The following theorem upper bounds $I(V;\mathcal{H}_{T})$.
@@
Let Assumption 3.1 hold. Under the assumptions and notations of Proposition 3.2, if $\Pi$ is a parametric decision space then
$I(V;\mathcal{H}_{T})\leq 2\bar{A}\epsilon^{2}T$ (5)
for $\epsilon\leq\epsilon_{0}$. If $\Pi$ is a non-parametric decision space then
$I(V;\mathcal{H}_{T})\leq\inf_{\delta>0}\left(\log\mathcal{N}\left(\rho,\Pi,\sqrt{\delta}\right)+T\delta\right).$ (6)
Now, substituting Theorem 3.3 in Proposition 3.2, and selecting $\Phi$ as we described above, results in the following Theorem.
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2\epsilon^{2}\bar{A}T+1}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}\right].$ (7)